### 25 Apr 2026 v1

# A proof of Bala’s general-$m$ representation of the harmonic numbers

Abstract.
For every nonzero integer $m$ and every integer $n\geq 1$, the $n$th harmonic number $H_{n}=1+\tfrac{1}{2}+\dots+\tfrac{1}{n}$ satisfies the identity
$H_{n}\;=\;\frac{1}{m}\,\sum_{k=1}^{n}\frac{(-1)^{k+1}}{k}\,\binom{mk}{k}\binom{n+(m-1)k}{n-k}$.
The cases $m=1$ and $m=2$ are classical; for general nonzero integer $m$ the identity was conjectured by P. Bala in the OEIS entry A001008 in 2022 and remained open. We prove it here, working throughout in $\mathbb{Q}[[x]]$. The proof reduces, via a substitution $u=x/(1-x)^{m}$, to two formal-power-series identities: a Lagrange–Bürmann evaluation of $\sum_{k\geq 1}\binom{mk}{k}u^{k}/k$, and the fixed-point fact that under that substitution the unique solution $v(u)$ of $v=u(1-v)^{m}$ is $v=x$. The argument extends verbatim to arbitrary complex $m\neq 0$.
Key words and phrases:
Harmonic number; binomial sum; Lagrange–Bürmann formula; formal power series; Fuss–Catalan; OEIS A001008
2020 Mathematics Subject Classification:
05A19, 05A10, 11B65, 11B83
1. Introduction
The numerators of the harmonic numbers
$H_{n}\;:=\;\sum_{i=1}^{n}\frac{1}{i},\qquad n\geq 1,\quad H_{0}:=0,$
form the integer sequence OEIS A001008. The On-Line Encyclopedia of Integer Sequences (OEIS) hosts a number of binomial-coefficient representations of $H_{n}$ contributed over time as comments to that entry. One of these, contributed by P. Bala on March 4, 2022, is the following parametric family.
Theorem 1 (Bala’s identity, conjectured 2022).
For every nonzero integer $m$ and every integer $n\geq 1$,
(1) $H_{n}\;=\;\frac{1}{m}\,\sum_{k=1}^{n}\frac{(-1)^{k+1}}{k}\,\binom{mk}{k}\binom{n+(m-1)k}{n-k}$.
[…]
$H_{n}=\tfrac{1}{2}\sum_{k=1}^{n}\frac{(-1)^{k+1}}{k}\binom{2k}{k}\binom{n+k}{n-k}$,
was contributed (also as a conjecture) to A001008 by G. Detlefs in April 2013 and proved shortly thereafter. For $m\geq 3$ and for every negative integer $m$, the identity (1) appears to have remained open until now.
In this note we prove Theorem 1 and observe that the proof in fact establishes the identity for every complex $m\neq 0$ when both sides are read as polynomials in $m$ for fixed $n$ (Remark 5).
The proof is via formal power series. In Section 2 we sum the right-hand side of (1) as an ordinary generating function $A_{m}(x)=\sum_{n}S_{m}(n)\,x^{n}$ and reduce it, under the substitution $u=x/(1-x)^{m}$, to a sum $T(u)=\sum_{k\geq 1}(-1)^{k+1}\binom{mk}{k}u^{k}/k$. Section 3 evaluates $T(u)$ via a Lagrange–Bürmann calculation. Section 4 closes the loop with a formal fixed-point lemma: under $u=x/(1-x)^{m}$, the unique solution $v(u)$ of $v=u(1-v)^{m}$ in $x\,\mathbb{Q}[[x]]$ is $v=x$. The conclusion $A_{m}(x)=-\log(1-x)/(1-x)=\sum_{n}H_{n}\,x^{n}$ is then immediate.
Throughout, we use the generalized binomial coefficient $\binom{a}{b}=a(a-1)\cdots(a-b+1)/b!$ for $a\in\mathbb{Z}$, $b\in\mathbb{Z}_{\geq 0}$, so that (1) is meaningful for all nonzero integer $m$. For negative $m$ this binomial may take negative or zero values term by term, but the sum still equals $H_{n}$ (see Example 7).
2. The generating function and a substitution
For nonzero integer $m$, write $S_{m}(n)$ for the right-hand side of (1), with the convention $S_{m}(0)=0$:
$S_{m}(n)\;:=\;\frac{1}{m}\,\sum_{k=1}^{n}\frac{(-1)^{k+1}}{k}\,\binom{mk}{k}\binom{n+(m-1)k}{n-k}$.
Let
$A_{m}(x)\;:=\;\sum_{n\geq 0}S_{m}(n)\,x^{n}\;\in\;\mathbb{Q}[[x]].$
Theorem 1 is equivalent to the formal power-series identity $A_{m}(x)=-\log(1-x)/(1-x)$, since $\log(1-x)/(1-x)=\sum_{n\geq 1}H_{n}\,x^{n}.$ We compute $A_{m}(x)$ in two steps.
Step 1: switch order of summation. We have
(2) $A_{m}(x)$ $=\frac{1}{m}\sum_{k\geq 1}\frac{(-1)^{k+1}}{k}\binom{mk}{k}\sum_{n\geq k}\binom{n+(m-1)k}{n-k}\,x^{n}.$
Setting j=nkj=n-k in the inner sum and using $\sum_{j\geq 0}\binom{j+r}{j}\,x^{j}=(1-x)^{-r-1}$ with $r=mk$,
(3) $\sum_{n\geq k}\binom{n+(m-1)k}{n-k}\,x^{n}=x^{k}\sum_{j\geq 0}\binom{j+mk}{j}\,x^{j}=\frac{x^{k}}{(1-x)^{mk+1}}.$
Step 2: introduce $u=x/(1-x)^{m}$. Substituting (3) into (2) and pulling out the factor $(1-x)^{-1}$,
(4) $A_{m}(x)=\frac{1}{m\,(1-x)}\,\sum_{k\geq 1}\frac{(-1)^{k+1}}{k}\binom{mk}{k}\,u^{k},\qquad u\;:=\;\frac{x}{(1-x)^{m}}.$
The proof of Theorem 1 therefore reduces to evaluating the inner sum in (4) as a formal power series in $u$, and then identifying the result after $u\mapsto x/(1-x)^{m}$.
3. A Lagrange–Bürmann identity
The next lemma is the key formal-series identity. Its proof uses the Lagrange–Bürmann formula in the coefficient-extraction form (see e.g. Stanley [3, Theorem 5.4.2]).
Lemma 2.
For every nonzero integer $m$,
(5) $\sum_{k\geq 1}\frac{1}{k}\binom{mk}{k}\,u^{k}\;=\;m\,\log(1+w),$
where $w=w(u)\in u\,\mathbb{Q}[[u]]$ is the unique formal power series satisfying
(6) $w\;=\;u\,(1+w)^{m}.$
Equivalently,
(7) $\sum_{k\geq 1}\frac{(-1)^{k+1}}{k}\binom{mk}{k}\,u^{k}\;=\;-\,m\,\log(1-v),$
where $v=v(u)\in u\,\mathbb{Q}[[u]]$ is the unique formal power series satisfying $v=u\,(1-v)^{m}$.
Proof.
The implication (5) $\Rightarrow$ (7) is obtained by replacing $u$ by u-u and setting v=w(u)v=-w(-u), which maps (6) to $v=u(1-v)^{m}$ and $\log(1+w(-u))$ to $\log(1-v)$.
So we prove (5). Equation (6) has a unique solution $w\in u\,\mathbb{Q}[[u]]$ because $\phi(w):=(1+w)^{m}$ has $\phi(0)=1\neq 0$ — even when $m$ is a negative integer, in which case $(1+w)^{m}$ is read as the formal series $\sum_{j\geq 0}\binom{-|m|}{j}\,w^{j}$. This is the standard Lagrange-inversion hypothesis.
Differentiate (6) implicitly with respect to $u$:
$w^{\prime}=(1+w)^{m}+u\,m\,(1+w)^{m-1}\,w^{\prime}.$
Solving for $w^{\prime}$ and using $u\,(1+w)^{m-1}=w/(1+w)$ (which follows from (6) after dividing both sides by $1+w$),
(8) $w^{\prime}\;=\;\frac{(1+w)^{m}}{1-mw/(1+w)}\;=\;\frac{(1+w)^{m+1}}{1+(1-m)\,w}.$
By the Lagrange–Bürmann coefficient-extraction formula ([3, Theorem 5.4.2]), for $\phi(w)=(1+w)^{m}$ and any formal series $f$,
$\sum_{k\geq 0}u^{k}\,[w^{k}]\bigl(f(w)\,\phi(w)^{k}\bigr)\;=\;\frac{f(w)}{1-u\,\phi^{\prime}(w)}.$
Applying this with $f(w)=1$ and using $w^{k}]\bigl((1+w)^{mk}\bigr)=\binom{mk}{k}$,
(9) $\sum_{k\geq 0}\binom{mk}{k}\,u^{k}\;=\;\frac{1}{1-u\,m\,(1+w)^{m-1}}\;=\;\frac{1}{1-mw/(1+w)}\;=\;\frac{1+w}{1+(1-m)\,w},$
where the second equality is again $u(1+w)^{m-1}=w/(1+w)$.
Subtracting the $k=0$ term and dividing by $u$,
(10) $\sum_{k\geq 1}\binom{mk}{k}\,u^{k-1}\;=\;\frac{1}{u}\!\left(\frac{1+w}{1+(1-m)w}-1\right)\;=\;\frac{1}{u}\cdot\frac{m\,w}{1+(1-m)\,w}\;=\;\frac{m\,(1+w)^{m}}{1+(1-m)\,w},$
where the last equality uses $1/u=(1+w)^{m}/w$ from (6).
Set
$T(u)\;:=\;\sum_{k\geq 1}\frac{1}{k}\binom{mk}{k}\,u^{k}\;\in\;u\,\mathbb{Q}[[u]].$
By (10), $dT/du$ equals the right-hand side of (10). Combining with (8) and using the chain rule with $w=w(u)$,
$\frac{dT}{dw}\;=\;\frac{dT}{du}\cdot\frac{du}{dw}\;=\;\frac{m\,(1+w)^{m}}{1+(1-m)\,w}\cdot\frac{1+(1-m)\,w}{(1+w)^{m+1}}\;=\;\frac{m}{1+w}.$
Integrating with $T(0)=0$ and $w(0)=0$ gives $T(u)=m\,\log(1+w)$, which is (5). ∎
Remark 3.
For each fixed $m$ the series $w$ is a Fuss–Catalan-type series: $1+w$ is the generating function of $m$-ary trees (shifted), and (1+w)/(1+(1m)w)(1+w)/(1+(1-m)w) in (9) is the well-known generating function for the binomials $\binom{mk}{k}$. For $m=2$, (6) solves to $1+w=2/(1+\sqrt{1-4u})$ and Lemma 2 specialises to the classical identity $\sum_{k\geq 1}\binom{2k}{k}u^{k}/k=2\log\bigl(2/(1+\sqrt{1-4u})\bigr)$.
4. The fixed-point substitution
Lemma 2 evaluates the inner sum in (4) once $v(u)$ is computed under $u=x/(1-x)^{m}$. The next lemma is the identification we need.
Lemma 4.
Let $v(u)\in u\,\mathbb{Q}[[u]]$ be the unique formal solution of $v=u(1-v)^{m}$. Under the formal substitution $u=x/(1-x)^{m}$, one has v(u)=xv(u)\;=\;x as formal power series in $x$.
Proof.
Let $V(x):=v\!\bigl(x/(1-x)^{m}\bigr)\in x\,\mathbb{Q}[[x]]$ be the formal-series composition. Substituting $u=x/(1-x)^{m}$ into $v=u(1-v)^{m}$ shows that $V$ satisfies
(11) $V(x)\;=\;\frac{x}{(1-x)^{m}}\,\bigl(1-V(x)\bigr)^{m}.$
Define the map $G:x\,\mathbb{Q}[[x]]\to x\,\mathbb{Q}[[x]]$ by
$G(\xi)\;:=\;\frac{x}{(1-x)^{m}}\,(1-\xi)^{m},$
where $(1-\xi)^{m}$ is the binomial formal series $\sum_{j\geq 0}\binom{m}{j}(-\xi)^{j}$ when $m\geq 0$, and the analogous expansion $\sum_{j\geq 0}\binom{-|m|+j-1}{j}(-1)^{j}\xi^{j}$ when $m<0$. In either case $(1-\xi)^{m}\in 1+\xi\,\mathbb{Q}[[\xi]]$ for $\xi(0)=0$. So $G$ is well-defined.
$G$ is a strict contraction in the $x$-adic topology. Indeed, if $\xi_{1}\equiv\xi_{2}\pmod{x^{N}}$ for some $N\geq 1$, then $(1-\xi_{1})^{m}-(1-\xi_{2})^{m}\in x^{N}\,\mathbb{Q}[[x]]$ by direct binomial expansion (each coefficient of the difference is a polynomial in $\xi_{1}-\xi_{2}$ with no constant term). Multiplying by $x/(1-x)^{m}\in x\,\mathbb{Q}[[x]]$ shifts the order by one further, so
$G(\xi_{1})\;\equiv\;G(\xi_{2})\pmod{x^{N+1}}$.
By the formal Banach fixed-point theorem (the $x$-adic completion of $\mathbb{Q}[x]$ is complete), $G$ has a unique fixed point in $x\,\mathbb{Q}[[x]]$.
Exhibiting the fixed point. Take $\xi=x$. Then $G(x)=x\,(1-x)^{m}/(1-x)^{m}=x$, so $x$ is a fixed point of $G$. By uniqueness, $V(x)=x$. ∎
5. Proof of Theorem 1
Proof of Theorem 1.
Substitute (7) into (4) with the $v$ in (7) viewed under $u=x/(1-x)^{m}$. By Lemma 4 this $v$ equals $x$, hence $\log(1-v)=\log(1-x)$ and
$A_{m}(x)\;=\;\frac{1}{m\,(1-x)}\cdot\bigl(-m\,\log(1-x)\bigr)\;=\;\frac{-\log(1-x)}{1-x}.$
The right-hand side is the standard generating function $\sum_{n\geq 1}H_{n}\,x^{n}$, obtained by Cauchy-multiplying the two series $\log(1-x)=\sum_{n\geq 1}x^{n}/n$ and $1/(1-x)=\sum_{n\geq 0}x^{n}$. Comparing coefficients of $x^{n}$ on both sides of $A_{m}(x)=-\log(1-x)/(1-x)$ yields $S_{m}(n)=H_{n}$ for every $n\geq 1$, which is (1). ∎
6. Remarks and examples
Remark 5 (Extension to complex $m$).
For each fixed $n\geq 1$, both sides of (1) are rational functions of $m$ (after clearing the leading $1/m$, they are polynomials in $m$ of degree at most n1n-1). The proof above establishes equality of these rational functions on the infinite set $\mathbb{Z}\setminus\{0\}$ of nonzero integers. Two rational functions agreeing on an infinite set are equal as rational functions, so (1) holds for every complex $m\neq 0$ when both sides are interpreted as rational functions of $m$.
Example 6 (Small $m$).
The cases $m=1,2,3,4$ of (1) read
$m=1:\quad$ $H_{n}=\sum_{k=1}^{n}\frac{(-1)^{k+1}}{k}\binom{n}{k}$,
$m=2:\quad$ $H_{n}=\tfrac{1}{2}\sum_{k=1}^{n}\frac{(-1)^{k+1}}{k}\binom{2k}{k}\binom{n+k}{n-k}$,
$m=3:\quad$ $H_{n}=\tfrac{1}{3}\sum_{k=1}^{n}\frac{(-1)^{k+1}}{k}\binom{3k}{k}\binom{n+2k}{n-k}$,
$m=4:\quad$ $H_{n}=\tfrac{1}{4}\sum_{k=1}^{n}\frac{(-1)^{k+1}}{k}\binom{4k}{k}\binom{n+3k}{n-k}$.
The first is the classical formula [1, (5.41)]; the second is Detlefs [6]; the rest are special cases of Theorem 1.
Example 7 (Negative $m$).
For m=1m=-1, the formula (1) becomes
$H_{n}\;=\;-\sum_{k=1}^{n}\frac{(-1)^{k+1}}{k}\,\binom{-k}{k}\binom{n-2k}{n-k}$,
where the binomials are the generalized binomials defined in Section 1. Most terms vanish because $\binom{n-2k}{n-k}=0$ for $\lceil n/2\rceil\leq k\leq n-1$ (the top index lies in $\{0,1,\dots,n-k-1\}$). For $n=4$ the only nonzero terms are $k=3$ and $k=4$, contributing $20/3$ and 35/4-35/4 respectively, with sum 25/12-25/12, so $(-25/12)=25/12=H_{4}$. The identity (1) has been numerically verified as exact rationals for $m\in\{-7,-5,-4,-3,-2,-1,1,2,\dots,8,10,15\}$ and $n\in\{1,\dots,40\}$ — all 640640 instances agree.
Remark 8 (Combinatorial interpretation).
The factor $\binom{mk}{k}/k$ has, up to a Fuss–Catalan factor of (m1)k+1(m-1)k+1, the meaning of the number of $m$-ary trees with $k$ internal nodes; cf. Stanley [3, §6.2] or Knuth [2, §7.2.1.6]. The factor $\binom{n+(m-1)k}{n-k}$ counts non-decreasing integer sequences of length nkn-k with values in $\{0,1,\dots,(m-1)k\}$. So (1) expresses $H_{n}$ as an inclusion–exclusion sum over a parameter $k$ that selects an $m$-ary tree size and a corresponding ”stretching” sequence. A purely bijective proof of (1) that does not pass through generating functions would be of independent interest.
Remark 9 (An infinite family of representations).
As $m$ ranges over $\mathbb{Z}\setminus\{0\}$, (1) gives an infinite family of distinct binomial-coefficient representations of $H_{n}$. By Remark 5, the family extends to a single algebraic relation in the parameter $m$. The ”right” coefficient identities — those obtained by extracting $m^{j}]$ from the polynomial identity in Remark 5 — encode binomial sums for the elementary symmetric polynomials in $\{1/1,1/2,\dots,1/n\}$ and may merit further investigation.
7. Numerical verification
The identity (1) and the intermediate identities (5), (7), and Lemma 4 have been verified by direct computation:
  • Both sides of (1) computed as exact rationals for $m\in\{-7,-5,-4,-3,-2,-1,1,2,3,4,5,6,7,8,10,15\}$ and $n\in\{1,2,\dots,40\}$ — all 640640 instances agree exactly.
  • The key Lagrange identity (5), its sign-alternating version (7), the classical evaluation $\sum_{k\geq 0}\binom{mk}{k}u^{k}=(1+w)/(1+(1-m)w)$, and the substitution $V(x)=x$ of Lemma 4 have all been checked as formal power-series identities to degree 88 in $u$ (resp. $x$) for $m\in\{-3,-2,-1,1,2,3,4,5\}$.
The verification scripts (Python; fractions.Fraction and sympy only) are available as supplementary material.
Acknowledgements
The author thanks Peter Bala for posing the conjecture proved here as a comment to OEIS sequence A001008, and the maintainers of the OEIS for hosting an indispensable repository of conjectured identities.