### 30 Apr 2026 v2

# A proof of Bala’s general-$m$ representation of the harmonic numbers

Abstract.
For every nonzero integer $m$ and every integer $n\geq 1$, the $n$th harmonic number $H_{n}=1+\tfrac{1}{2}+\dots+\tfrac{1}{n}$ satisfies the identity
$H_{n}\;=\;\frac{1}{m}\,\sum_{k=1}^{n}\frac{(-1)^{k+1}}{k}\,\binom{mk}{k}\binom{n+(m-1)k}{n-k}$.
The cases $m=1$ and $m=2$ are classical; for general nonzero integer $m$ the identity was conjectured by P. Bala in the OEIS entry A001008 in 2022 and remained open. We prove it here, working throughout in $\mathbb{Q}[[x]]$. The proof reduces, via a substitution $u=x/(1-x)^{m}$, to two formal-power-series identities: a Lagrange–Bürmann evaluation of $\sum_{k\geq 1}\binom{mk}{k}u^{k}/k$, and the fixed-point fact that under that substitution the unique solution $v(u)$ of $v=u(1-v)^{m}$ is $v=x$. The argument extends verbatim to arbitrary complex $m\neq 0$.
Key words and phrases:
Harmonic number; binomial sum; Lagrange–Bürmann formula; formal power series; Fuss–Catalan; OEIS A001008
2020 Mathematics Subject Classification:
05A19, 05A10, 11B65, 11B83
1. Introduction
The numerators of the harmonic numbers
$H_{n}\;:=\;\sum_{i=1}^{n}\frac{1}{i},\qquad n\geq 1,\quad H_{0}:=0,$
form the integer sequence OEIS A001008. The On-Line Encyclopedia of Integer Sequences (OEIS) hosts a number of binomial-coefficient representations of $H_{n}$ contributed over time as comments to that entry. One of these, contributed by P. Bala on March 4, 2022, is the following parametric family.
Theorem 1 (Bala’s identity, conjectured 2022).
For every nonzero integer $m$ and every integer $n\geq 1$,
(1) $H_{n}\;=\;\frac{1}{m}\,\sum_{k=1}^{n}\frac{(-1)^{k+1}}{k}\,\binom{mk}{k}\binom{n+(m-1)k}{n-k}$.
[…]
$H_{n}=\tfrac{1}{2}\sum_{k=1}^{n}\frac{(-1)^{k+1}}{k}\binom{2k}{k}\binom{n+k}{n-k}$,
was contributed (also as a conjecture) to A001008 by G. Detlefs in April 2013 and proved shortly thereafter. For $m\geq 3$ and for every negative integer $m$, the identity (1) appears to have remained open until now.
In this note we prove Theorem 1 and observe that the proof in fact establishes the identity for every complex $m\neq 0$ when both sides are read as polynomials in $m$ for fixed $n$ (Remark 5).
The proof is via formal power series. In Section 2 we sum the right-hand side of (1) as an ordinary generating function $A_{m}(x)=\sum_{n}S_{m}(n)\,x^{n}$ and reduce it, under the substitution $u=x/(1-x)^{m}$, to a sum $T(u)=\sum_{k\geq 1}(-1)^{k+1}\binom{mk}{k}u^{k}/k$. Section 3 evaluates $T(u)$ via a Lagrange–Bürmann calculation. Section 4 closes the loop with a formal fixed-point lemma: under $u=x/(1-x)^{m}$, the unique solution $v(u)$ of $v=u(1-v)^{m}$ in $x\,\mathbb{Q}[[x]]$ is $v=x$. The conclusion $A_{m}(x)=-\log(1-x)/(1-x)=\sum_{n}H_{n}\,x^{n}$ is then immediate.
Throughout, we use the generalized binomial coefficient $\binom{a}{b}=a(a-1)\cdots(a-b+1)/b!$ for $a\in\mathbb{Z}$, $b\in\mathbb{Z}_{\geq 0}$, so that (1) is meaningful for all nonzero integer $m$. For negative $m$ this binomial may take negative or zero values term by term, but the sum still equals $H_{n}$ (see Example 7).
2. The generating function and a substitution
For nonzero integer $m$, write $S_{m}(n)$ for the right-hand side of (1), with the convention $S_{m}(0)=0$:
$S_{m}(n)\;:=\;\frac{1}{m}\,\sum_{k=1}^{n}\frac{(-1)^{k+1}}{k}\,\binom{mk}{k}\binom{n+(m-1)k}{n-k}$.
Let
$A_{m}(x)\;:=\;\sum_{n\geq 0}S_{m}(n)\,x^{n}\;\in\;\mathbb{Q}[[x]].$
Theorem 1 is equivalent to the formal power-series identity $A_{m}(x)=-\log(1-x)/(1-x)$, since $\log(1-x)/(1-x)=\sum_{n\geq 1}H_{n}\,x^{n}.$ We compute $A_{m}(x)$ in two steps.
Step 1: switch order of summation. We have
(2) $A_{m}(x)$ $=\frac{1}{m}\sum_{k\geq 1}\frac{(-1)^{k+1}}{k}\binom{mk}{k}\sum_{n\geq k}\binom{n+(m-1)k}{n-k}\,x^{n}.$
Setting j=nkj=n-k in the inner sum and using $\sum_{j\geq 0}\binom{j+r}{j}\,x^{j}=(1-x)^{-r-1}$ with $r=mk$,
(3) $\sum_{n\geq k}\binom{n+(m-1)k}{n-k}\,x^{n}=x^{k}\sum_{j\geq 0}\binom{j+mk}{j}\,x^{j}=\frac{x^{k}}{(1-x)^{mk+1}}.$
Step 2: introduce $u=x/(1-x)^{m}$. Substituting (3) into (2) and pulling out the factor $(1-x)^{-1}$,
(4) $A_{m}(x)=\frac{1}{m\,(1-x)}\,\sum_{k\geq 1}\frac{(-1)^{k+1}}{k}\binom{mk}{k}\,u^{k},\qquad u\;:=\;\frac{x}{(1-x)^{m}}.$
The proof of Theorem 1 therefore reduces to evaluating the inner sum in (4) as a formal power series in $u$, and then identifying the result after $u\mapsto x/(1-x)^{m}$.
3. A Lagrange–Bürmann identity
The next lemma is the key formal-series identity. Its proof uses the Lagrange–Bürmann formula in the coefficient-extraction form (see e.g. Stanley [3, Theorem 5.4.2]; for a short elementary inductive proof of the Lagrange inversion formula, see Surya and Warnke [4]).
Lemma 2.
For every nonzero integer $m$,
(5) $\sum_{k\geq 1}\frac{1}{k}\binom{mk}{k}\,u^{k}\;=\;m\,\log(1+w),$
where $w=w(u)\in u\,\mathbb{Q}[[u]]$ is the unique formal power series satisfying
(6) $w\;=\;u\,(1+w)^{m}.$
Equivalently,
(7) $\sum_{k\geq 1}\frac{(-1)^{k+1}}{k}\binom{mk}{k}\,u^{k}\;=\;-\,m\,\log(1-v),$
where $v=v(u)\in u\,\mathbb{Q}[[u]]$ is the unique formal power series satisfying $v=u\,(1-v)^{m}$.
Proof.
The implication (5) $\Rightarrow$ (7) is obtained by replacing $u$ by u-u and setting v=w(u)v=-w(-u), which maps (6) to $v=u(1-v)^{m}$ and $\log(1+w(-u))$ to $\log(1-v)$.
So we prove (5). Equation (6) has a unique solution $w\in u\,\mathbb{Q}[[u]]$ because $\phi(w):=(1+w)^{m}$ has $\phi(0)=1\neq 0$ — even when $m$ is a negative integer, in which case $(1+w)^{m}$ is read as the formal series $\sum_{j\geq 0}\binom{-|m|}{j}\,w^{j}$. This is the standard Lagrange-inversion hypothesis.
Differentiate (6) implicitly with respect to $u$:
$w^{\prime}=(1+w)^{m}+u\,m\,(1+w)^{m-1}\,w^{\prime}.$
Solving for $w^{\prime}$ and using $u\,(1+w)^{m-1}=w/(1+w)$ (which follows from (6) after dividing both sides by $1+w$),
(8) $w^{\prime}\;=\;\frac{(1+w)^{m}}{1-mw/(1+w)}\;=\;\frac{(1+w)^{m+1}}{1+(1-m)\,w}.$
By the Lagrange–Bürmann coefficient-extraction formula ([3, Theorem 5.4.2]), for $\phi(w)=(1+w)^{m}$ and any formal series $f$,
$\sum_{k\geq 0}u^{k}\,[w^{k}]\bigl(f(w)\,\phi(w)^{k}\bigr)\;=\;\frac{f(w)}{1-u\,\phi^{\prime}(w)}.$
Applying this with $f(w)=1$ and using $w^{k}]\bigl((1+w)^{mk}\bigr)=\binom{mk}{k}$,
(9) $\sum_{k\geq 0}\binom{mk}{k}\,u^{k}\;=\;\frac{1}{1-u\,m\,(1+w)^{m-1}}\;=\;\frac{1}{1-mw/(1+w)}\;=\;\frac{1+w}{1+(1-m)\,w},$
where the second equality is again $u(1+w)^{m-1}=w/(1+w)$.
Subtracting the $k=0$ term and dividing by $u$,
(10) $\sum_{k\geq 1}\binom{mk}{k}\,u^{k-1}\;=\;\frac{1}{u}\!\left(\frac{1+w}{1+(1-m)w}-1\right)\;=\;\frac{1}{u}\cdot\frac{m\,w}{1+(1-m)\,w}\;=\;\frac{m\,(1+w)^{m}}{1+(1-m)\,w},$
where the last equality uses $1/u=(1+w)^{m}/w$ from (6).
Set
$T(u)\;:=\;\sum_{k\geq 1}\frac{1}{k}\binom{mk}{k}\,u^{k}\;\in\;u\,\mathbb{Q}[[u]].$
By (10), $dT/du$ equals the right-hand side of (10). Combining with (8) and using the chain rule with $w=w(u)$,
$\frac{dT}{dw}\;=\;\frac{dT}{du}\cdot\frac{du}{dw}\;=\;\frac{m\,(1+w)^{m}}{1+(1-m)\,w}\cdot\frac{1+(1-m)\,w}{(1+w)^{m+1}}\;=\;\frac{m}{1+w}.$
Integrating with $T(0)=0$ and $w(0)=0$ gives $T(u)=m\,\log(1+w)$, which is (5). ∎
Remark 3.
For each fixed $m$ the series $w$ is a Fuss–Catalan-type series: $1+w$ is the generating function of $m$-ary trees (shifted), and (1+w)/(1+(1m)w)(1+w)/(1+(1-m)w) in (9) is the well-known generating function for the binomials $\binom{mk}{k}$. For $m=2$, (6) solves to $1+w=2/(1+\sqrt{1-4u})$ and Lemma 2 specialises to the classical identity $\sum_{k\geq 1}\binom{2k}{k}u^{k}/k=2\log\bigl(2/(1+\sqrt{1-4u})\bigr)$.
4. The fixed-point substitution
Lemma 2 evaluates the inner sum in (4) once $v(u)$ is computed under $u=x/(1-x)^{m}$. The next lemma is the identification we need.
Lemma 4.
Let $v(u)\in u\,\mathbb{Q}[[u]]$ be the unique formal solution of $v=u(1-v)^{m}$. Under the formal substitution $u=x/(1-x)^{m}$, one has v(u)=xv(u)\;=\;x as formal power series in $x$.
Proof.
Let $V(x):=v\!\bigl(x/(1-x)^{m}\bigr)\in x\,\mathbb{Q}[[x]]$ be the formal-series composition. Substituting $u=x/(1-x)^{m}$ into $v=u(1-v)^{m}$ shows that $V$ satisfies
(11) $V(x)\;=\;\frac{x}{(1-x)^{m}}\,\bigl(1-V(x)\bigr)^{m}.$
Define the map $G:x\,\mathbb{Q}[[x]]\to x\,\mathbb{Q}[[x]]$ by
$G(\xi)\;:=\;\frac{x}{(1-x)^{m}}\,(1-\xi)^{m},$
where $(1-\xi)^{m}$ is the binomial formal series $\sum_{j\geq 0}\binom{m}{j}(-\xi)^{j}$ when $m\geq 0$, and the analogous expansion $\sum_{j\geq 0}\binom{-|m|+j-1}{j}(-1)^{j}\xi^{j}$ when $m<0$. In either case $(1-\xi)^{m}\in 1+\xi\,\mathbb{Q}[[\xi]]$ for $\xi(0)=0$. So $G$ is well-defined.
$G$ is a strict contraction in the $x$-adic topology. Indeed, if $\xi_{1}\equiv\xi_{2}\pmod{x^{N}}$ for some $N\geq 1$, then $(1-\xi_{1})^{m}-(1-\xi_{2})^{m}\in x^{N}\,\mathbb{Q}[[x]]$ by direct binomial expansion (each coefficient of the difference is a polynomial in $\xi_{1}-\xi_{2}$ with no constant term). Multiplying by $x/(1-x)^{m}\in x\,\mathbb{Q}[[x]]$ shifts the order by one further, so
$G(\xi_{1})\;\equiv\;G(\xi_{2})\pmod{x^{N+1}}$.
By the formal Banach fixed-point theorem (the $x$-adic completion of $\mathbb{Q}[x]$ is complete), $G$ has a unique fixed point in $x\,\mathbb{Q}[[x]]$.
Exhibiting the fixed point. Take $\xi=x$. Then $G(x)=x\,(1-x)^{m}/(1-x)^{m}=x$, so $x$ is a fixed point of $G$. By uniqueness, $V(x)=x$. ∎
5. Proof of Theorem 1
Proof of Theorem 1.
Substitute (7) into (4) with the $v$ in (7) viewed under $u=x/(1-x)^{m}$. By Lemma 4 this $v$ equals $x$, hence $\log(1-v)=\log(1-x)$ and
$A_{m}(x)\;=\;\frac{1}{m\,(1-x)}\cdot\bigl(-m\,\log(1-x)\bigr)\;=\;\frac{-\log(1-x)}{1-x}.$
The right-hand side is the standard generating function $\sum_{n\geq 1}H_{n}\,x^{n}$, obtained by Cauchy-multiplying the two series $\log(1-x)=\sum_{n\geq 1}x^{n}/n$ and $1/(1-x)=\sum_{n\geq 0}x^{n}$. Comparing coefficients of $x^{n}$ on both sides of $A_{m}(x)=-\log(1-x)/(1-x)$ yields $S_{m}(n)=H_{n}$ for every $n\geq 1$, which is (1). ∎
6. Remarks and examples
Remark 5 (Extension to complex $m$).
For each fixed $n\geq 1$, both sides of (1) are rational functions of $m$ (after clearing the leading $1/m$, they are polynomials in $m$ of degree at most n1n-1). The proof above establishes equality of these rational functions on the infinite set $\mathbb{Z}\setminus\{0\}$ of nonzero integers. Two rational functions agreeing on an infinite set are equal as rational functions, so (1) holds for every complex $m\neq 0$ when both sides are interpreted as rational functions of $m$.
Example 6 (Small $m$).
The cases $m=1,2,3,4$ of (1) read
$m=1:\quad$ $H_{n}=\sum_{k=1}^{n}\frac{(-1)^{k+1}}{k}\binom{n}{k}$,
$m=2:\quad$ $H_{n}=\tfrac{1}{2}\sum_{k=1}^{n}\frac{(-1)^{k+1}}{k}\binom{2k}{k}\binom{n+k}{n-k}$,
$m=3:\quad$ $H_{n}=\tfrac{1}{3}\sum_{k=1}^{n}\frac{(-1)^{k+1}}{k}\binom{3k}{k}\binom{n+2k}{n-k}$,
$m=4:\quad$ $H_{n}=\tfrac{1}{4}\sum_{k=1}^{n}\frac{(-1)^{k+1}}{k}\binom{4k}{k}\binom{n+3k}{n-k}$.
The first is the classical formula [1, (5.41)]; the second is Detlefs [7]; the rest are special cases of Theorem 1.
Example 7 (Negative $m$).
For m=1m=-1, the formula (1) becomes
$H_{n}\;=\;-\sum_{k=1}^{n}\frac{(-1)^{k+1}}{k}\,\binom{-k}{k}\binom{n-2k}{n-k}$,
where the binomials are the generalized binomials defined in Section 1. Most terms vanish because $\binom{n-2k}{n-k}=0$ for $\lceil n/2\rceil\leq k\leq n-1$ (the top index lies in $\{0,1,\dots,n-k-1\}$). For $n=4$ the only nonzero terms are $k=3$ and $k=4$, contributing $20/3$ and 35/4-35/4 respectively, with sum 25/12-25/12, so $(-25/12)=25/12=H_{4}$. The identity (1) has been numerically verified as exact rationals for $m\in\{-7,-5,-4,-3,-2,-1,1,2,\dots,8,10,15\}$ and $n\in\{1,\dots,40\}$ — all $640$ instances agree.
Remark 8 (Combinatorial interpretation).
The factor $\binom{mk}{k}/k$ has, up to a Fuss–Catalan factor of (m1)k+1(m-1)k+1, the meaning of the number of $m$-ary trees with $k$ internal nodes; cf. Stanley [3, §6.2] or Knuth [2, §7.2.1.6]. The factor $\binom{n+(m-1)k}{n-k}$ counts non-decreasing integer sequences of length nkn-k with values in $\{0,1,\dots,(m-1)k\}$. So (1) expresses $H_{n}$ as an inclusion–exclusion sum over a parameter $k$ that selects an $m$-ary tree size and a corresponding "stretching" sequence. A purely bijective proof of (1) that does not pass through generating functions would be of independent interest.
Remark 9 (An infinite family of representations).
As $m$ ranges over $\mathbb{Z}\setminus\{0\}$, (1) gives an infinite family of distinct binomial-coefficient representations of $H_{n}$. By Remark 5, the family extends to a single algebraic relation in the parameter $m$. The "right" coefficient identities — those obtained by extracting $m^{j}]$ from the polynomial identity in Remark 5 — encode binomial sums for the elementary symmetric polynomials in $\{1/1,1/2,\dots,1/n\}$ and may merit further investigation.
7. Numerical verification
The identity (1) and the four intermediate identities used in the proof — (5), (7), the classical evaluation (9), and the substitution $V(x)=x$ of Lemma 4 — have been verified by direct computation:
  • Both sides of (1) computed as exact rationals (Python fractions.Fraction) for $m\in\{-7,-5,-4,-3,-2,-1,1,2,3,4,5,6,7,8,10,15\}$ and $n\in\{1,2,\dots,40\}$ — all $640$ instances agree exactly. See Appendix A.
  • The four intermediate formal-power-series identities have all been checked coefficient-by-coefficient up to degree $8$ in $u$ (resp. $x$) for $m\in\{-3,-2,-1,1,2,3,4,5\}$. See Appendix B.
The two verification scripts are self-contained Python (standard library only, no third-party dependencies) and reproduce in seconds. They are provided in full in Appendices A and B.
Acknowledgements
The author thanks Peter Bala for posing the conjecture proved here as a comment to OEIS sequence A001008, and the maintainers of the OEIS for hosting an indispensable repository of conjectured identities. The author is also grateful to Lutz Warnke for kindly drawing attention to the recent elementary inductive proof of the Lagrange inversion formula in [4].
Appendix A Exact-rational verification of identity (1)
The script below evaluates both sides of identity (1) as exact rationals (fractions.Fraction), without any floating point, for every pair $(m,n)$ with $m\in\{-7,-5,-4,-3,-2,-1,1,2,3,4,5,6,7,8,10,15\}$ and $n\in\{1,\dots,40\}$640640 instances. All instances match.
Listing 1: verify_bala.py: exact-rational verification of (1).
1"""
2Verify Peter Bala's conjecture (2022) on A001008:
3For any nonzero integer m,
[…]
5
6where H(n) = 1 + 1/2 + ... + 1/n.
7
8Strategy: compute both sides as exact Fraction, compare equality.
9"""
10
11from fractions import Fraction
12from math import comb
13
14
15def harmonic(n: int) -> Fraction:
[…]
17 h = Fraction(0)
18 for i in range(1, n + 1):
19 h += Fraction(1, i)
20 return h
21
22
23def gen_binom(top: int, k: int) -> int:
24 """Compute generalized binomial coefficient C(top, k) for integer top, k >= 0.
25 Uses falling factorial: top * (top-1) * ... * (top - k + 1) / k!.
26 Equals math.comb when top >= 0, and uses the extension when top < 0.
27 """
28 if k < 0:
29 return 0
30 if top >= 0:
31 if top < k:
32 return 0
33 return comb(top, k)
34 # top < 0: use falling factorial. C(top, k) = (-1)^k * C(-top + k - 1, k)
35 return (-1) ** k * comb(-top + k - 1, k)
36
37
38def bala_rhs(n: int, m: int) -> Fraction:
39 """Compute Peter Bala's conjectured formula RHS:
[…]
41 """
42 s = Fraction(0)
43 for k in range(1, n + 1):
44 sign = 1 if k % 2 == 1 else -1
45 term = Fraction(sign, k) * gen_binom(m * k, k) * gen_binom(n + (m - 1) * k, n - k)
46 s += term
47 return s / m
48
49
50def main() -> None:
[…]
52 print()
53
54 # m must be nonzero integer. Negative m uses generalized (falling-factorial) binomial.
55 test_ms = [-7, -5, -4, -3, -2, -1, 1, 2, 3, 4, 5, 6, 7, 8, 10, 15]
56 test_ns = list(range(1, 41))
57
58 all_ok = True
59 for m in test_ms:
60 for n in test_ns:
61 lhs = harmonic(n)
62 rhs = bala_rhs(n, m)
63 ok = lhs == rhs
64 if not ok:
65 print(f" m={m:+d} n={n:2d} H(n)={lhs} RHS={rhs} [FAIL]")
66 all_ok = False
67 print(f" m={m:+d}: all n in 1..{test_ns[-1]} verified OK")
68 print()
69 print("All tests passed." if all_ok else "FAILURES detected.")
70
71
72if __name__ == "__main__":
73 main()
Output (abridged):
[…]
                                          * C(mk,k) * C(n+(m-1)k, n-k)
  m=-7: all n in 1..40 verified  OK
  m=-5: all n in 1..40 verified  OK
  m=-4: all n in 1..40 verified  OK
  m=-3: all n in 1..40 verified  OK
  m=-2: all n in 1..40 verified  OK
  m=-1: all n in 1..40 verified  OK
  m=+1: all n in 1..40 verified  OK
  m=+2: all n in 1..40 verified  OK
  m=+3: all n in 1..40 verified  OK
  m=+4: all n in 1..40 verified  OK
  m=+5: all n in 1..40 verified  OK
  m=+6: all n in 1..40 verified  OK
  m=+7: all n in 1..40 verified  OK
  m=+8: all n in 1..40 verified  OK
  m=+10: all n in 1..40 verified  OK
  m=+15: all n in 1..40 verified  OK
All tests passed.
The script depends only on the Python standard library (fractions.Fraction and math.comb). The total runtime is a fraction of a second on commodity hardware. Generalised binomial coefficients $\binom{a}{b}$ for negative $a$ are evaluated by the falling-factorial extension $\binom{-k}{j}=(-1)^{j}\binom{k+j-1}{j}$, matching the convention of Section 1.
Appendix B Formal-power-series verification of the four intermediate identities
This second script verifies, by truncating to degree $N=8$ and comparing coefficients, the four formal-power-series identities used in the proof of Theorem 1:
  1. (i)
    the Lagrange identity (5) of Lemma 2,
  2. (ii)
    its sign-alternating version (7),
  3. (iii)
    the classical evaluation $\sum_{k\geq 0}\binom{mk}{k}u^{k}=(1+w)/(1+(1-m)w)$ of (9), and
  4. (iv)
    the substitution $V(x)=x$ of Lemma 4.
For each $m\in\{-3,-2,-1,1,2,3,4,5\}$, the implicit fixed-point equation $w=u(1+w)^{m}$ is solved iteratively as a truncated power series, and the four identities are then verified coefficient-by-coefficient. All $4\times 8=32$ instances pass.
Listing 2: check_proof.py: formal-power-series verification of the four intermediate identities.
1"""
2Sanity-check the key identities used in the proof of Bala's harmonic-number conjecture:
3
[…]
6 where v = u (1-v)^m.
7
8Both checked as power-series identities, by truncating to degree N and comparing
9coefficients of [u^k].
10
11We also verify the key substitution: if u = x/(1-x)^m, then the power-series
12solution v(u) of v = u(1-v)^m satisfies v = x, and so
13
[…]
15
16confirming Bala's identity S_m(n) = H_n.
17"""
18
19from fractions import Fraction
20from math import comb
21
22
23def gen_binom(top: int, k: int) -> int:
24 """Generalized binomial C(top, k) for integer top and k >= 0.
25 Uses falling factorial extension for negative top."""
26 if k < 0:
27 return 0
28 if top >= 0:
29 if top < k:
30 return 0
31 return comb(top, k)
32 return (-1) ** k * comb(-top + k - 1, k)
33
34
35def truncate_log1p(series_coeffs: list[Fraction], N: int) -> list[Fraction]:
36 """Given coefficients of f(u) = sum c_k u^k with c_0 = 0,
37 compute coefficients of log(1 + f(u)) up to degree N.
38 """
39 # Use formal power series logarithm via integration of f'/(1+f).
40 # Easier: log(1+f) where f = sum c_k u^k, c_0 = 0.
41 # Method: g = log(1+f) satisfies (1+f) g' = f' so we can extract coefs recursively:
[…]
49 f = list(series_coeffs)
50 while len(f) <= N:
51 f.append(Fraction(0))
52 d = [Fraction(0)] * (N + 1)
53 for n in range(1, N + 1):
54 s = Fraction(0)
55 for i in range(1, n):
56 s += f[i] * (n - i) * d[n - i]
57 d[n] = f[n] - s / n
58 return d
59
60
61def lagrange_solve(phi: list[Fraction], N: int) -> list[Fraction]:
62 """Given phi = [phi_0, phi_1, ...] coefficients of a power series with phi_0 != 0,
[…]
64 truncated to degree N.
65
66 Uses the recursion implied by w = u * phi(w): write w = u * phi(w) and
67 repeatedly substitute. We use the formal power series approach.
68 """
[…]
70 w = [Fraction(0)] * (N + 1)
71 for _ in range(N + 1):
72 # phi(w) up to degree N (composition).
73 phi_w = [Fraction(0)] * (N + 1)
74 # phi_w = sum_j phi_j * w^j
75 wpow = [Fraction(0)] * (N + 1)
76 wpow[0] = Fraction(1) # w^0 = 1
77 for j in range(len(phi)):
78 for d in range(N + 1):
79 phi_w[d] += phi[j] * wpow[d]
80 # Update wpow = wpow * w (truncated)
81 new_wpow = [Fraction(0)] * (N + 1)
82 for a in range(N + 1):
83 if wpow[a] == 0:
84 continue
85 for b in range(N + 1 - a):
86 new_wpow[a + b] += wpow[a] * w[b]
87 wpow = new_wpow
88 # New w = u * phi_w, i.e., shift right by 1.
89 new_w = [Fraction(0)] * (N + 1)
90 for d in range(N):
91 new_w[d + 1] = phi_w[d]
92 w = new_w
93 return w
94
95
96def series_mul(a: list[Fraction], b: list[Fraction], N: int) -> list[Fraction]:
97 out = [Fraction(0)] * (N + 1)
98 for i in range(min(len(a), N + 1)):
99 if a[i] == 0:
100 continue
101 for j in range(min(len(b), N + 1 - i)):
102 out[i + j] += a[i] * b[j]
103 return out
104
105
106def series_inverse(a: list[Fraction], N: int) -> list[Fraction]:
107 """Multiplicative inverse of a power series with a[0] != 0."""
108 inv = [Fraction(0)] * (N + 1)
109 inv[0] = 1 / a[0]
110 for n in range(1, N + 1):
111 s = Fraction(0)
112 for i in range(1, min(n, len(a) - 1) + 1):
113 s += a[i] * inv[n - i]
114 inv[n] = -s / a[0]
115 return inv
116
117
118def main() -> None:
119 N = 8
120 print(f"Verifying identities to degree {N}.")
121 print()
122
123 for m in (-3, -2, -1, 1, 2, 3, 4, 5):
[…]
125 lhs = [Fraction(0)] * (N + 1)
126 for k in range(1, N + 1):
127 lhs[k] = Fraction(1, k) * gen_binom(m * k, k)
128
129 # Compute w with phi(w) = (1+w)^m (coefficients), truncated to deg N.
130 # For m >= 0, phi is polynomial of degree m. For m < 0, it's an infinite series.
131 phi_len = (m + 1) if m >= 0 else (N + 1)
132 phi = [Fraction(gen_binom(m, j)) for j in range(phi_len)]
133 w = lagrange_solve(phi, N)
134
135 # RHS: m * log(1 + w).
136 log1pw = truncate_log1p(w, N)
137 rhs = [m * c for c in log1pw]
138
139 ok = lhs == rhs
140 print(f" m={m} Identity (I) [sum (1/k) C(mk,k) u^k = m log(1+w)] OK={ok}")
141 if not ok:
142 print(f" LHS: {lhs}")
143 print(f" RHS: {rhs}")
144
145 # Check identity (II): sum (-1)^(k+1)/k * C(mk,k) u^k = -m log(1-v) with v = u(1-v)^m.
146 lhs2 = [Fraction(0)] * (N + 1)
147 for k in range(1, N + 1):
148 lhs2[k] = Fraction((-1) ** (k + 1), k) * gen_binom(m * k, k)
149 # phi2(v) = (1-v)^m, truncated to deg N for m < 0.
150 phi2_len = (m + 1) if m >= 0 else (N + 1)
151 phi2 = [Fraction((-1) ** j * gen_binom(m, j)) for j in range(phi2_len)]
152 v = lagrange_solve(phi2, N)
153 # log(1 - v) = log(1 + (-v)).
154 neg_v = [-c for c in v]
155 log1mv = truncate_log1p(neg_v, N)
156 rhs2 = [-m * c for c in log1mv]
157 ok2 = lhs2 == rhs2
158 print(f" m={m} Identity (II) [sum (-1)^(k+1)/k C(mk,k) u^k = -m log(1-v)] OK={ok2}")
159 if not ok2:
160 print(f" LHS: {lhs2}")
161 print(f" RHS: {rhs2}")
162
163 # Check that when u = x/(1-x)^m, v = x.
164 # This is a substitution check at the formal level: use power series in x.
165 Nx = N
166 u_of_x = [Fraction(0)] * (Nx + 1)
[…]
168 for j in range(Nx):
169 u_of_x[j + 1] = Fraction(gen_binom(m + j - 1, j))
170 # Compute v_of_x via composition v(u(x)).
171 # We need v as a power series in u, then substitute u = u_of_x.
172 # v(u) = sum v_k u^k where v from above.
173 # Compose: v_of_x = sum v_k * (u_of_x)^k.
174 v_of_x = [Fraction(0)] * (Nx + 1)
175 upowx = [Fraction(0)] * (Nx + 1)
176 upowx[0] = Fraction(1)
177 for k in range(Nx + 1):
178 for d in range(Nx + 1):
179 v_of_x[d] += v[k] * upowx[d]
180 new_upowx = [Fraction(0)] * (Nx + 1)
181 for a in range(Nx + 1):
182 if upowx[a] == 0:
183 continue
184 for b in range(Nx + 1 - a):
185 new_upowx[a + b] += upowx[a] * u_of_x[b]
186 upowx = new_upowx
187 target = [Fraction(0)] * (Nx + 1)
188 target[1] = Fraction(1) # v = x
189 ok3 = v_of_x == target
190 print(f" m={m} Substitution u = x/(1-x)^m gives v = x OK={ok3}")
191 if not ok3:
192 print(f" v_of_x: {v_of_x}")
193
[…]
195 bcoef = [Fraction(gen_binom(m * k, k)) for k in range(N + 1)]
196 one_plus_w = [Fraction(0)] * (N + 1)
197 one_plus_w[0] = Fraction(1)
198 for k in range(N + 1):
199 one_plus_w[k] += w[k]
200 one_plus_1mw = [Fraction(0)] * (N + 1)
201 one_plus_1mw[0] = Fraction(1)
202 for k in range(N + 1):
203 one_plus_1mw[k] += (1 - m) * w[k]
204 denom_inv = series_inverse(one_plus_1mw, N)
205 rhs5 = series_mul(one_plus_w, denom_inv, N)
206 ok5 = bcoef == rhs5
207 print(f" m={m} Identity (5) [sum C(mk,k) u^k = (1+w)/(1+(1-m)w)] OK={ok5}")
208 if not ok5:
209 print(f" LHS: {bcoef}")
210 print(f" RHS: {rhs5}")
211 print()
212
213
214if __name__ == "__main__":
215 main()
Output (abridged):
Verifying identities to degree 8.
  m=-3  Identity (I)  [sum (1/k) C(mk,k) u^k = m log(1+w)]    OK=True
  m=-3  Identity (II) [sum (-1)^(k+1)/k ...     = -m log(1-v)] OK=True
  m=-3  Substitution u = x/(1-x)^m gives v = x                 OK=True
  m=-3  Identity (5)  [sum C(mk,k) u^k = (1+w)/(1+(1-m)w)]     OK=True
  m=+5  Identity (I)                                           OK=True
  m=+5  Identity (II)                                          OK=True
  m=+5  Substitution u = x/(1-x)^m gives v = x                 OK=True
  m=+5  Identity (5)                                           OK=True
The script depends only on the Python standard library. Power-series multiplication, multiplicative inverse, formal logarithm, and the Lagrange-type fixed-point solver ($w=u\,\phi(w)$ for $\phi$ a polynomial or formal series) are all implemented from scratch in $\sim 100$ lines, so the verification itself has no third-party dependencies and is straightforward to audit.